OpenStudy (anonymous):

Find the equation of the line tangent to f(x)=x/((x^2)+3) at (0,0)

OpenStudy (moonlitfate):

Just be sure, the function given is: \[f(x) = \frac{ x }{ x^2+3 }\]

OpenStudy (moonlitfate):

And this is for Calculus?

OpenStudy (moonlitfate):

If it is, then the first thing that you are going to do is take the derivative of the function using the quotient rule.

OpenStudy (moonlitfate):

@jazzie -- do you know how to do that?

OpenStudy (anonymous):

Yes this is for calculus ! I am preparing for a test and am kinda lost when it comes to this @MoonlitFate

OpenStudy (moonlitfate):

Ah, okay. Well, does taking the derivative of that confuse you?

OpenStudy (anonymous):

Isnt the derivative (3-x^2)/((x^2) +3)^2)

OpenStudy (moonlitfate):

Yes. Sorry, I was working it out.

OpenStudy (moonlitfate):

Okay, I'm used to doing simpler problems, but what I would usually do is set x to 0 in the derivative and work it out to get the y-intercept.

OpenStudy (anonymous):

Okay if it helps these are my options x – 3y = 0 y – 3x = 0 x + 3y = 0 y + 3x = 0 Does not exist.

OpenStudy (moonlitfate):

I got does not exist; but, I'll verify it, real quick.

OpenStudy (moonlitfate):

Hmm, I did that wrong...

OpenStudy (anonymous):

Okay the tangent is where the two lines touch ?

OpenStudy (moonlitfate):

The tangent line is a line that touches a curve at a certain, but it doesn't cross over it.

OpenStudy (anonymous):

they just touch. okay so i could graph it and find the point ?